OpenStudy (anonymous):

when calculating final velocity im having a hard time with units? how do i know what units to convert and what units to leave as they are?

OpenStudy (anonymous):

velocity = m/s

OpenStudy (anonymous):

I have that I just don't know what unit my answer should be in

OpenStudy (anonymous):

if you are calculating a velocity you need the units to say m/s

OpenStudy (anonymous):

and acceleration is in m/s^2

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so much help.... procrastinators dream

OpenStudy (anonymous):

one more question.... would I convert everything in the problem to the unit my answer is in?

OpenStudy (anonymous):

well throughout your work units should carry or cancel. use the units to check your answer

OpenStudy (anonymous):

acceleration: 4m/s^2 vi:0km/h vf: 965 km/h find the distance

OpenStudy (plainntall):

You will have to cancel out the hours and kilometers and replace them with meters and seconds. 0 km/h is just 0 m/s, zero is zero 965km/h 965km/h(1000m/km) ; you cancel out the km by dividing by it and putting its meter equivalent on top 965000m/h 965000m/h(1h/3600 s) ; you cancel out the hour by multiplying by 1 hour and dividing by the equivalent seconds of 3600s vf = 268 m/s

OpenStudy (anonymous):

ok maybe im more confused than I thought... if I do that my answer comes out in meters per second but I need it to be in km?? and when I convert my answer back to km its different than the correct answer given by my teacher!

OpenStudy (anonymous):

u see in an equation all the physical quantities need ot be in the same units... doesn't matter which system u use for example.. u are using equation of motion which involves time, velocity and displacement if ur velocity is in km/hr.. u need to put time in hr, and displacement in Km.. if u have acceleration u will have to put it in km/hr^2.. which no one does.. and therefore we usually convert everything to SI.. that is meters and seconds.. but its not a necessity

OpenStudy (plainntall):

Oh, sorry, so you need the acceleration in km/h^2?

OpenStudy (anonymous):

but if I convert everything to meters and seconds and do the equation my answer is still wrong....

OpenStudy (anonymous):

car went to maximum speed of 965km/h. started at rest and underwent constant accel with a magnitude of 4.0m/s^2. what distance would she have had to travel in order to reach the maximum speed. just walk through this..

OpenStudy (plainntall):

If it is the distance then convert acceleration to m/hr^2 which is 4m/s^2 4m/s^2(1km/1000m)(3600^2s^2/1h^2) 51840 km/h^2 So that gives us d=(Vf^2-vi^2)/2a =((965km/h)^2-(0km/h)^2)/(2(51840km/h^2) d=8.98 km

OpenStudy (anonymous):

ok it worked on that one but not the next........ :( thanks for ur help the rider travels with an initial speed of 8.0m/s before speeding up. the rider then travels 40.0 m at a constant acceleration of 2m/s^2 what is the riders final speed?

OpenStudy (anonymous):

I don't get how it comes out wrong? all these units are already where they need to be so should I be able to just plug and chug?

OpenStudy (plainntall):

I think so, use vf^2=vi^2+2ad

OpenStudy (plainntall):

That should be vf^2=(8)^2m^2/s^2+2(2m/s^2)(40.0m)

OpenStudy (anonymous):

ok last one.. I think I have the hang of it somewhat! but the skid marks of a jet were 9.6 km long. if the jets acceleration was -2m/s^2 what was the cars initial velocity?

OpenStudy (anonymous):

how would u do this? u cant convert km to m/s so idk what to do here

OpenStudy (plainntall):

Vi^2=Vf^2-2ad where vf=0 vi^2=-2ad You don't convert to m/s it is a distance not a speed. Vi^2=-2(-2m/s^2)(9.6km)(1000m/1km)

OpenStudy (plainntall):

the units comput to m^2/s^2) which when you take the square root are m/s which is right fof speed

OpenStudy (anonymous):

you literally saved me 3 hours... idk if you know how thankful I am!! thank you sooooooooo much!

OpenStudy (plainntall):

glad I could help

OpenStudy (anonymous):

like honestly go buy urself something nice